OpenStudy (anonymous):

I am trying to think of some tricky linear or quadratic equation problems (the "solve for x" problem sets, no imaginary roots please) for a group of minors. Have any ideas?

OpenStudy (nenadmatematika):

X^2-5X+6=0 :D

OpenStudy (anonymous):

That is too easy

OpenStudy (turingtest):

work backwards, what do you want the roots to be? whatever a and b they are you can make the equation from \[(x-a)(x-b)=0\]

OpenStudy (turingtest):

classics also include things with hidden quadratics, like\[ax^{2n}+bx^n+c=0\]which is quadratic in \(x^n\) or\[an^{2x}+bn^x+c=0\]which is quadratic in \(n^x\)